OpenStudy (anonymous):

3/4(x-12)=12 I need some help trying to find what is the value of x also 3/4y-12=12 my answer for this one is 16 am I right or wrong?

pooja195 (pooja195):

Hi :) Welcome to openstudy! 3/4(x-12)=12 Start by distributing 3/4 

pooja195 (pooja195):

We know that 3/4 * x=3/4x 3/4 * -12=?

OpenStudy (anonymous):

okay so we need to divide the numerator by the number in the parenthesis

OpenStudy (anonymous):

or do we multiply

pooja195 (pooja195):

Or you can \[\huge~\rm~ \frac{ 3 \times -12 }{ 4 \times 1 }=? \]

pooja195 (pooja195):

3 * -12=?

OpenStudy (dayakar):

@pooja195 ,i think you are mistaken\[\frac{ 3 }{ 4}(x-12) =12\]

pooja195 (pooja195):

?

OpenStudy (anonymous):

yeah like that pooja

pooja195 (pooja195):

That's what i have

pooja195 (pooja195):